TrailPacker produces rugged backpacks for outdoor sports (hiking, rock climbing, etc.) Their backpacks are sold at many specialty outdoor stores across the country. The cost of manufacturing and marketing their backpacks, at their normal factory volume of 20,000 backpacks per month, is shown in the table below. TrailPacker sells these backpacks for $50 each. TrailPacker is making a small profit, but they would prefer to increase their Operating Income.

Per Unit

Per Unit

Unit Manufacturing Cost

$ 7.00

Variable Materials

$ 11.00

Variable Labor

$ 6.00

Variable Overhead

$ 10.00

Total Unit Manufacturing Cost

$ 34.00

Unit Marketing Costs

Variable Marketing Costs

$ 2.00

Fixed Marketing Costs

$ 4.00

Total Unit Marketing Costs

$ 6.00

TrailPacker is thinking of increasing sales by offering backpacks with aluminum frames. The investment needed for adding aluminum frames to their manufacturing process would increase fixed overhead costs by $50,000 per month. The variable materials cost (only variable material costs – not all variable costs) would increase by $15.00 per backpack. Market research estimates that the aluminum frame backpacks would sell for $70 each, and volume would increase 10%.

A) Prepare a revised monthly Contribution Margin Income Statement to include the revenues, detailed costs and income if TrailPacker chooses to manufacture and sell backpacks with aluminum frames instead of their regular backpacks.

B) What is the new break-even point in units for the aluminum frame backpacks?

C) What is the new break-even point in sales dollars for the aluminum frame backpacks?

D) If volume did not increase when making the aluminum frame backpacks (stayed the same as original monthly volume), is TrailPacker better off producing aluminum frame backpacks or their basic backpacks? Support your answer with data in contribution margin income statement format.
Answer #1

A. Contribution margin income statement

Sales revenue 20,000*1.1*70 =$1,540,000

Less: variable costs

Manufacturing cost =22000*7 =$154000

Variable material 22000*26 =$572,000

Variable labor 22000*6 =$132000

Variable overhead 22000*10 =$220,000

Variable marketing =22000*2 =$44,000

Total variable cost =$1,122,000

Less: fixed costs 50,000+4*20,000 =$130,000

Net operating income =$992,000

B.break even point = fixed costs/contribution margin per unit

=130,000/19

=6842.11 Units

C.in sales = 130,000/27.14%

=$478,997.79

D. At original volume

Sales revenue =20,000*70 =$1400,000

Less variable costs 20,000*51 =$1020000

Less: fixed costs =$130,000

Net operating income =$250,000

Profit in case of original backpacks =20,000*14 - 80,000 =$200,000

Hence, better off with aluminium frame backpacks